# Evolutionary Optimization of Physics-Informed Neural Networks: Evo-PINN Frontiers and Opportunities

### v3

@@ III Optimization and Generalization Challenges
However, PINN’s advantages come at the cost of increased difficulty in optimizing PINN model parameters, with many early papers showing long training times and limited predictive performance on more complex problems [31, 41, 42, 43, 39, 40]. Moreover, parameterized PINNs can still face issues with generalizability since the accuracy of interpolation-based predictions on a new problem is not guaranteed, and training a well-generalizable PINN model may not be an easy task. In this section, we shed light on some contributing factors to the optimization and generalization challenges—2 pre-dominant hurdles to the wide-spread adoption of PINNs in real-world applications today—and how they may originate from fundamental limitations in gradient-based approaches.
@@ III-A Challenges in Optimizing PINN Models
In many studies involving PINN, it was observed that the training iterations required for convergence far exceed those needed for typical data-driven DNN. Unsuccessful PINN training (premature convergence) can lead to unacceptable (unphysical) simulation outcomes. We relate the optimization difficulty of PINNs to a few factors below.
@@ III-A1 Complex Loss Landscape
The physics-informed loss usually consists of multiple PDE and IC/BC loss terms (Equation 3). The different loss terms can display distinct scales and characteristics during PINN training. For instance, the scale of PDE loss can span several orders of magnitude depending on the underlying differential operators and the complexity of the solution, and this range is only revealed during training. Additionally, some loss terms change more dramatically than others during training, as shown by prior analyses through the lens of neural tangent kernel (NTK) [82, 83, 84]. These characteristics make the proper scaling of different loss terms challenging.
In addition, the PINN optimization landscapes are littered with spurious minima and are more complex than those of usual data-driven DNN models. Fig. 5 illustrates the loss landscapes of PINN and DNN models for the convection-diffusion problem presented in Subsection II-B. To investigate the topological differences in their loss landscapes, we first computed the Hessians of their respective loss functions (i.e., $\mathcal{L}_{\text{PINN}}$ or $\mathcal{L}_{\text{DNN}}$) with respect to the model parameters: $\mathbf{H}=\nabla^{2}\ \mathcal{L}(\boldsymbol{\boldsymbol{w}})$. We then extracted the leading two eigenvectors and eigenvalues of the Hessian (i.e., first 2 principal Hessian directions) which correspond to the directions of highest curvature in the loss landscape. By perturbing the trained models along these dominant eigenvectors and evaluating the resulting loss values, we can obtain a more informative characterization of the local topology than perturbing the model parameters in random directions [39]. This visualization provides evidence of a much higher degree of complexity and ruggedness of the loss landscape for PINN compared to usual data-driven DNN models. The smoothness of the latter is conducive to gradient-based algorithms that can easily find descent paths from the point of initialization to a good local optimum that fits the training data well. This outcome corroborates the primacy of SGD and its variants in today’s deep learning tasks. In contrast, the contradictory descent directions from different loss terms encountered during PINN training can adversely impact optimization [85, 86]. It is worth nothing that when the loss terms are purely physics-based (e.g., PDE and IC/BC), they are not theoretically contradictory. However, during the optimization process, especially when the solution is still far from the true optimum, the local gradient directions of these loss components can still become practically conflicting. This issue arises due to the combined effects of neural network expressivity, the structure of the loss landscape (e.g., scales and convergence dynamics), and the limitations of optimizers, which can cause certain objectives to dominate or suppress others during early or intermediate stages of training.
@@ III-A2 Initial Local Minimum Trap and Spectral Bias
Another closely related issue stems from the known saturation property of the activation functions commonly used for PINN. Such saturation property can cause their backpropagation gradients to vanish whenever the network weights are shifted towards a larger distribution, as driven by the need to approximate higher frequencies and steep gradients common in physics. This makes further gradient descent of the network weights difficult, adversely causing a learning bias towards smooth functions (lower frequencies are learned better than higher frequencies). This known tendency—also called spectral bias—has been studied through the lens of NTK [92, 93], revealing a significant discrepancy in convergence rates for different frequency components in the output during training. While this has been regarded as a natural regularizer with potential benefits to the generalization performance of DNNs, the consequences are highly deleterious for PINNs. Spectral bias makes it difficult for PINNs to accurately learn complex nonlinear physics behaviors [94], e.g., in the vicinity of steep gradients characterizing shock waves. These local discrepancies further deter propagation of the correct initial and boundary conditions into the computation domain, which has been suggested to be critical for successful PINN training [88, 95].
@@ III-A3 Sensitivity to Sample Size and Location
The training samples of the PINN model are typically uniformly drawn from the input domain and are equally important everywhere due to their physical correspondence with the spatio-temporal extent of the problem. While the use of mini-batch samples is a standard procedure in gradient-based training, it remains an open question how one can efficiently reduce the batch size (and hence the training cost) while maintaining the desired PINN model accuracy [98]. Using a large batch size is crucial for the accuracy of PINNs due to the potential need to learn high frequencies or complex structures in the a priori unknown target function [99]. However, this inevitably increases training costs. Furthermore, oversampling on a high loss region may adversely slow convergence since it can be harder to learn such local complex gradient behaviour—a cause of stiffness during training [56, 57, 42, 43]. On the other hand, an overly small batch size will result in noisy (big jump in loss values) and inconsistent local gradient signals [100], thereby causing issues during gradient descent.
@@ III-A4 Current Research Directions
Since the seminal work of Raissi et al. [31], the architecture design and training methods for PINNs have continued to evolve. Recent advancements often combine multiple learning strategies to tackle the aforementioned PINN optimization issues [101, 102, 103]. Curriculum learning improves convergence by initiating PINNs training on a better-conditioned, simpler version of the original problem and then systematically advancing to target problem [39, 41]. Physics-informed learning often leads to premature convergence when dealing with challenging problems, however, the same PDE equation with an altered parameters or IC/BCs may become a simpler problem. In this scenario, one can start training the PINN model with the easier problem and then gradually vary the problem during training until reaching the target problem. This approach assumes that after each transition step, the training will bring the model close enough to the optimal solution of the next (harder) problem, providing a good starting point for the next round of training. Hence, curriculum learning can be viewed as an initialization process to guide the model toward the desired solution utilizing a priori understanding of how the system behaves as its problem parameters change. Adaptive loss-balancing strategy can also be considered as a form of curriculum learning, where the training is initiated with the loss weights corresponding to a simpler optimization problem. It involves assigning appropriate weights to balance different terms in the PINN loss functions [104, 86, 83, 50, 82, 84, 85], and is regarded as an important technique to improve convergence.
@@
Recent studies have proposed enhancements to PINN learning capability by leveraging network architectures that incorporate Fourier feature embeddings, skip connections, and random weight factorization [101, 102, 103]. PINNs with skip connections and periodic activation functions retain a greater proportion of gradients during backpropagation from output to input, helping to mitigate spectral bias caused by activation saturation [105, 89, 91]. Adaptive techniques that adjust the slope of activation functions via trainable multiplicative factors have also been introduced to further control saturation and enhance training efficiency [106, 107]. These methods enable PINNs to embed inductive biases—such as high-frequency behavior characteristic of certain PDEs—directly into the network architecture. Additionally, physics-informed cell representations have been developed to address spectral bias away from the input origin by disentangling network parameters from the input coordinates [108]. Convolutional architectures also achieve this parameter-input decoupling, allowing for larger and more expressive networks that may scale more effectively to complex, large-scale problems [61].
@@
Some studies have proposed using numerical differentiation (ND)-based methods for the differential operators required by the physics-informed loss, which is much cheaper to compute as compared to automatic differentiation-evaluated loss [56, 57, 62]. More importantly, ND are more robust to sample density and potentially advantageous in the context of reducing sample requirement in establishing a strong correlation between physics-informed loss and model accuracy. This is because ND methods, by definition, use Taylor series expansion (TSE) to construct a low-order local approximation for the derivative terms. They can provide structural biases and discourage extreme gradients to reduce overfitting and make PINNs easier to learn the approximate solution [56]. They can also better connect neighboring samples and facilitate the propagation of information from the IC/BC to interior sample points, thereby reducing any tendency to overfit especially near boundaries [57].
@@ III-B Challenges in Learning Generalizable PINN Model
Given both simulation (hypothesis evaluation) and inference capabilities, PINNs are versatile and can be applied to various applications common in science such as model-predictive-control, scenario planning, and design exploration [109, 110, 111, 112]; see the example in Fig. 6. The commonality to these applications is the need for multiple accurate and physics-compliant predictions across a series of related problems. These repeated evaluations usually differ by modifications to the physics scenario such as changes in geometric shape, PDE parameter, IC and BC.
III-B1 How Generalizable Is A PINN Model?
@@
Most parameterized PINN models (ref. to Subsection II-C; they are designed to learn a parametric family of physics scenarios) make predictions on new scenarios unseen during training through interpolation. Contrary to its promise, the interpolation is usually physics-agnostic, with similar negative implications for generalizability as the predictions from typical DNNs [55]. Given the difficulties in optimizing a PINN model for a single physics scenario, the cost of performing physics-informed learning across a set of scenarios can be prohibitive [113]. The effectiveness of these parameterized PINNs for prediction of new scenarios then depends on the complexity of the problem. Moreover, the problem’s parameterization needs to be determined at the learning stage, which reduces the flexibility of the parameterized PINN model when applied to new physics scenarios.
III-B2 How to Train A Generalizable PINN Model in The Data-Scarce Regime?
@@
Meta-learning approaches provide a systematic framework for discovering good weight initializations that can be quickly adapted—with minimal training—to arrive at physically precise predictions in every new scenario (task) for a given problem distribution. Meta-learning of a generic PINN model initialization can also allow generalization across diverse, multi-modal problem distributions, hence it holds immense promise for addressing generalization challenge of PINN [114, 115, 116, 117, 113, 118, 119, 120, 121].
@@
There are still a number of unresolved issues in achieving a generalizable PINN model via meta-learning. A fundamental assumption of popular gradient-based meta-learning algorithms—such as model-agnostic meta-learning (MAML) algorithm [122] and its variant, Reptile [123]—is that the models can learn well on each task-specific training after a few gradient descent updates from the initialization. However, this is usually not the case for PINNs because physics-informed losses are much more challenging to optimize. They require more extensive training to achieve good convergence [113]. With only a handful of task-specific gradient descent updates, gradient-based meta-learning algorithms may still be prone to getting stuck in local minima as the gradients collected from a batch of diverse training tasks can be noisy or even deceptive. As a result, recent PINN studies have shown inferior results for MAML, potentially due to insufficient task-specific adaptation and non-convergence in the meta-learning algorithm’s outer loop optimization [118, 119].
IV Evo-PINN: Physics-Informed Learning with Evolutionary Algorithms
@@
Despite their theoretical advantage in diverse scientific domains, Section III highlighted several practical challenges in today’s PINN learning and optimization. The cost of optimization, navigating spectral bias and local minima traps, has even seeded a sense of pessimism among computational scientists about the utility of these techniques over standard numerical methods [124, 125]. In this section, we present our hypothesis that many algorithmic ideas developed in the computational intelligence community may in fact be uniquely well positioned to tackle the issues plaguing PINNs. One exciting research direction is to achieve physics-informed learning by synergizing evolutionary algorithms with local gradient descent signals, thereby leveraging the best of both worlds to quickly arrive at physically accurate and generalizable PINN models. In what follows, we examine this and other promising algorithmic directions that we believe warrant sustained investigation. These directions highlight key opportunities for the evolutionary optimization of PINNs (Evo-PINNs) to advance scientific machine learning.
@@ IV-A1 Neuroevolution Can Avoid Local Minima Traps
The operators of EAs are typically based on principles of natural selection. They do not require derivative computations of the loss functions being optimized and are therefore gradient-free. Importantly, as population-based methods, the implicit parallelism of EAs implies that they have lesser tendency to get trapped in a single local minimum [127, 52]. What’s more, since their search is typically invariant to order-preserving transformations of the loss function [128], EAs are less susceptible to flat output functions with small (vanishing) gradients. For these reasons, they are deemed as potential alternatives to SGD, particularly in the context of scientific machine learning where the goal is to converge to physically precise solutions represented by global optima with zero PINN loss [51]. In addition to its global search capacity, an evolved population naturally provides an ensemble of neural networks that affords the possibility of accurate predictions with robust uncertainty quantification—as was shown through the application of a swarm-based evolutionary procedure to PINN optimization [129, 130]. The ability to provide uncertainty quantification and probabilistic forecasts is crucial for many scientific modelling and control problems [131].
@@ IV-A2 Pareto Optimization Balances Physics and Data Losses
The existence of multiple loss terms—see Eqs. (2) and (3)—makes PINN training intrinsically a multi-objective optimization problem. Since the terms $\mathcal{L}_{data},\mathcal{L}_{pde},\mathcal{L}_{ic}$ and $\mathcal{L}_{bc}$ represent different physical quantities, their magnitudes may be incommensurable. In this case, the role of the scalarizing weights is not only to express preferences between the losses, but also to compensate for differences in their magnitudes. It is worth noting that a perfect physics solution would, in theory, simultaneously satisfy all PINN loss terms. However, in practice, the expressive capacity of a neural network is always limited by its (necessarily) finite size. As argued in [97, 132], capacity limitations naturally give rise to conflicts and trade-offs between loss terms. This, in conjunction with the extreme ruggedness of the physics-informed learning objective, implies that a neural network, despite being a theoretically universal function approximator, is not guaranteed to find a solution that simultaneously minimizes all the terms [49]. On the other hand, when data-driven terms are included, such as observational data or empirical measurements, true contradictions may arise. This can occur when the data is noisy, incomplete, or governed by partially unknown physics that differs from the modeled equations. In such cases, the multi-objective optimization must reconcile inherently conflicting objectives between data fitting and physics adherence. Identifying an appropriate aggregation of objectives to arrive at good (physically satisfactory) trade-off solution is also non-trivial [133] Moreover, the shape of the Pareto front formed by the multiple objectives is not known beforehand [134], with simple weighted-sum aggregations being insufficient to support all parts of non-convex fronts [135].
@@
Multi-objective EAs (MOEAs) are tailor-made to address the issues mentioned above [136]. In particular, Pareto dominance-based variants of MOEAs [137] do not require the multiple loss terms to be scalarized, and hence are less sensitive to incommensurable objectives. Their population-based search mechanisms also allow for free-form convergence towards the entire Pareto front in a single run of the algorithm, irrespective of the front’s shape. These salient features have naturally begun to attract the attention of PINN researchers, although associated developments are still at a nascent stage. In [138], the non-dominated sorting genetic algorithm [139] was hybridized with a variant of SGD (namely, Adam [140]) for training the parameters of a PINN. The exploration capacity of the EA was shown to help overcome local minima and enable precise satisfaction of the initial and boundary conditions encoded in the physics-informed loss. Given the computational cost of a hybrid framework, a pure population-based multi-objective search algorithm was studied for PINN training in [130], leading to calibrated uncertainty quantification by examining the empirical distribution of the ensemble’s predictions.
@@
In addition to using MOEAs for PINN parameter training, there have been efforts towards evolving the model’s hyperparameters (including the scalarizing weights and activation functions). The method yields a Pareto set of hyperparameters where each setting offers an optimal trade-off between the data and physics losses. Practitioners can then decide post-hoc which setting they wish to pursue [141]. An analysis of such post-hoc decision-making revealed that solutions close to the convex bulge of the Pareto front typically led to performant PINNs [133]. Evolutionary algorithmists will immediately notice a connection between this observation and the literature on knee-point driven MOEAs [142], which could serve as a valuable resource for hyperparameter tuning in PINNs.

### v5

@@ III Optimization and Generalization Challenges
However, PINN’s advantages come at the cost of increased difficulty in optimizing PINN model parameters, with many early papers showing long training times and limited predictive performance on more complex problems [31, 39, 40, 38]. Unsuccessful PINN training (premature convergence) can lead to unacceptable (unphysical) simulation outcomes. Moreover, parameterized PINNs can still face issues with generalizability since the accuracy of interpolation-based predictions on a new problem is not guaranteed, and training a well-generalizable PINN model may not be an easy task. In this section, we shed light on some contributing factors to the optimization and generalization challenges—two pre-dominant hurdles to the wide-spread adoption of PINNs in real-world applications today—and how they may originate from fundamental limitations in gradient-based approaches.
@@ III-A Challenges in Optimizing PINN Models
In many studies involving PINN, it was observed that the training iterations required for convergence far exceed those needed for typical data-driven DNN. We relate the optimization difficulty of PINNs to a few factors below.
@@ III-A1 Complex Loss Landscape
The physics-informed loss usually consists of multiple PDE and IC/BC loss terms (Equation 3), each with distinct scales and characteristics during PINN training. For instance, the scale of PDE loss can span several orders of magnitude depending on the underlying differential operators and the complexity of the solution, and this range is only revealed during training. Additionally, some loss terms change more dramatically than others during training, as shown by prior analyses through the lens of neural tangent kernel (NTK) [75, 76]. These characteristics make the proper scaling of different loss terms challenging.
In addition, the PINN optimization landscapes are littered with spurious minima and are more complex than those of usual data-driven DNN models. Fig. 5 illustrates the loss landscapes of PINN and DNN models for the convection-diffusion problem presented in Subsection II-B. To investigate the topological differences in their loss landscapes, we first computed the Hessians of their respective loss functions (i.e., $\mathcal{L}_{\text{PINN}}$ or $\mathcal{L}_{\text{DNN}}$) with respect to the model parameters: $\mathbf{H}=\nabla^{2}\ \mathcal{L}(\boldsymbol{\boldsymbol{w}})$. We then extracted the leading two eigenvectors and eigenvalues of the Hessian (i.e., first two principal Hessian directions) which correspond to the directions of highest curvature in the loss landscape. By perturbing the trained models along these dominant eigenvectors and evaluating the resulting loss values, we can obtain a more informative characterization of the local topology than perturbing the model parameters in random directions [38]. This visualization provides evidence of a much higher degree of complexity and ruggedness of the loss landscape for PINN compared to usual data-driven DNN models. The smoothness of the latter is conducive to gradient-based algorithms that can easily find descent paths from the point of initialization to a good local optimum that fits the training data well. This outcome corroborates the primacy of SGD and its variants in today’s deep learning tasks. In contrast, the contradictory descent directions from different loss terms encountered during PINN training can adversely impact optimization [77, 78]. It is worth noting that when the loss terms are purely physics-based (e.g., PDE and IC/BC), they are not theoretically contradictory. However, during the optimization process, especially when the solution is still far from the true optimum, the local gradient directions of these loss components can still become practically conflicting. This issue arises due to the combined effects of neural network expressivity, the structure of the loss landscape (e.g., scales and convergence dynamics), and the limitations of optimizers, which can cause certain objectives to dominate or suppress others during early or intermediate stages of training.
@@ III-A2 Initial Local Minimum Trap and Spectral Bias
Another closely related issue stems from the known saturation property of activation functions commonly used for PINN. Such saturation property can cause backpropagated gradients to vanish whenever the network weights are shifted towards a larger distribution, as driven by the need to approximate higher frequencies and steep gradients common in physics. This makes further gradient descent of the network weights difficult, adversely causing a learning bias towards smooth functions (lower frequencies are learned better than higher frequencies). This known tendency—also called spectral bias—has been studied through the lens of NTK [84, 85], revealing a significant discrepancy in convergence rates for different frequency components in the output during training. While this has been regarded as a natural regularizer with potential benefits to the generalization performance of DNNs, the consequences are highly deleterious for PINNs. Spectral bias makes it difficult for PINNs to accurately learn complex nonlinear physics behaviors [86], e.g., in the vicinity of steep gradients characterizing shock waves. These local discrepancies further deter propagation of the correct initial and boundary conditions into the computation domain, which has been suggested to be critical for successful PINN training [80, 87].
@@ III-A3 Sensitivity to Sample Size and Location
The training samples of the PINN model are typically uniformly drawn from the input domain and are equally important everywhere due to their physical correspondence with the spatio-temporal extent of the problem. While the use of mini-batch samples is a standard procedure in gradient-based training, it remains an open question how one can efficiently reduce the batch size (and hence the training cost) while maintaining the desired PINN model accuracy [90]. Using a large batch size is crucial for the accuracy of PINNs due to the potential need to learn high frequencies or complex structures in the a priori unknown target function [91]. However, this inevitably increases training costs. Furthermore, oversampling on a high loss region may adversely slow convergence since it can be harder to learn such local complex gradient behaviour—a cause of stiffness during training [54, 39, 40]. On the other hand, an overly small batch size will result in noisy (big jump in loss values) and inconsistent local gradient signals [92], thereby causing issues during gradient descent.
@@ III-A4 Current Research Directions
Since the seminal work of Raissi et al. [31], architecture design and training methods for PINNs have continued to evolve. Recent advancements often combine multiple learning strategies to tackle the aforementioned PINN optimization issues [93, 94, 95]. Curriculum learning improves convergence by initiating training on a better-conditioned, simpler version of the original problem and then systematically advancing to target problem [38]. Physics-informed learning often succumbs to premature convergence for challenging problems, however, the same PDE equation with altered parameters or IC/BCs may be a simpler problem. In this scenario, one can start training the PINN model on the easier problem and then gradually vary the problem during training to approach the target problem. This approach assumes that each training step brings the model closer to the optimal solution of the next (harder) problem, thereby improving the next round of training. Hence, curriculum learning can be viewed as an initialization process to guide the model toward the desired solution utilizing a priori understanding of how the system behaves as its problem parameters change. Adaptive loss-balancing strategy can also be considered as a form of curriculum learning, where the training is initiated with the loss weights corresponding to a simpler optimization problem. It involves assigning appropriate weights to balance different terms in the PINN loss functions [78, 76, 47, 75, 77], and is regarded as an important technique to improve convergence.
@@
Recent studies have proposed enhancements to PINN learning by leveraging network architectures that incorporate Fourier feature embeddings, skip connections, and random weight factorization [93, 94, 95]. PINNs with skip connections and periodic activation functions retain a greater proportion of gradients during backpropagation from output to input, thereby mitigating spectral bias caused by activation saturation [96, 81, 83]. Adaptive techniques that adjust the slope of activation functions have also been introduced to further control saturation and enhance training efficiency [97, 98]. These methods embed inductive biases—such as high-frequency behavior characteristic of certain PDEs—directly into the network architecture. Additionally, physics-informed cell representations have been developed to address spectral bias away from the input origin by disentangling network parameters from the input coordinates [99]. Convolutional architectures also achieve this parameter-input decoupling, allowing for larger, more expressive networks that scale effectively to complex, large-scale problems [58].
@@
Some studies have proposed using numerical differentiation (ND)-based methods for the differential operators in the physics-informed loss, with ND being potentially cheaper to compute relative to automatic differentiation [53, 54]. More importantly, ND is more robust to sample density and potentially advantageous in the context of reducing sample requirement in establishing a strong correlation between physics-informed loss and model accuracy. This is because ND methods, by definition, use Taylor series expansion (TSE) to construct a low-order local approximation for the derivative terms. They can provide structural biases and discourage extreme gradients to reduce overfitting to better learn the approximate solution [53]. They can also connect neighboring samples and facilitate the propagation of information from the IC/BC to interior sample points, thereby reducing any tendency to overfit especially near boundaries [54].
@@ III-B Challenges in Learning Generalizable PINN Model
Given both simulation (hypothesis evaluation) and inference capabilities, PINNs are versatile and can be applied to various applications common in science such as model-predictive-control, scenario planning, and design exploration [100, 101, 102, 103]; see Fig. 6. The commonality to these applications is the need for multiple accurate and physics-compliant predictions across a series of related problems. These repeated evaluations usually differ by modifications to the physics scenario such as changes in geometric shape, PDE parameter, IC, and BC.
@@
1) How Generalizable Is A PINN Model? Most parameterized PINN models (ref. to Subsection II-C; they are designed to learn a parametric family of physics scenarios) make predictions on new scenarios unseen during training through interpolation. This interpolation is usually physics-agnostic, with similar negative implications for generalizability as predictions from typical DNNs [52]. Consequently, the effectiveness of parameterized PINNs depends significantly on problem complexity. Given the difficulties in optimizing a PINN model for a single scenario, the cost of performing physics-informed learning across a set of scenarios can be prohibitive [104]. Moreover, the problem’s parameterization needs to be determined at the learning stage, which reduces the flexibility of the parameterized PINN model when applied to new physics scenarios.
@@
2) How to Train A Generalizable PINN Model in The Data-Scarce Regime? Meta-learning approaches provide a systematic framework for discovering good weight initializations that can be quickly adapted—with minimal training—to new scenarios (tasks) in a problem distribution. Hence, meta-learning of a generic PINN model initialization can similarly accelerate physically precise predictions across diverse, multi-modal problem distributions, and holds immense promise for addressing generalization challenge of PINN [105, 106, 107, 108, 104, 109, 110, 111, 112].
@@
There are still a number of unresolved issues in meta-learning of a generalizable PINN model. Recent PINN studies showed inferior results for model-agnostic meta-learning (MAML), potentially due to insufficient task-specific adaptation and non-convergence in the meta-learning algorithm’s outer loop optimization [109, 110]. A fundamental assumption of popular gradient-based meta-learning algorithms—such as MAML [113] and its variant, Reptile [114]—is that the models can learn well on each task after a few gradient descent updates from the initialization. However, this is not the case for PINNs because physics-informed losses are much more challenging to optimize and require extensive training [104]. With only a handful of task-specific gradient descent updates, gradient-based meta-learning algorithms may remain susceptible to local minima from noisy or even deceptive gradients.
IV Evo-PINN: Physics-Informed Learning with
Evolutionary Algorithms
@@ IV Evo-PINN: Physics-Informed Learning with
Despite their theoretical advantage in diverse scientific domains, Section III has highlighted several practical challenges in today’s PINN learning and optimization. The cost of optimization, navigating spectral bias and local minima traps, has even seeded a sense of pessimism among computational scientists about the utility of these techniques over standard numerical methods [115, 116]. In this section, we present our hypothesis that many algorithmic ideas developed in the computational intelligence community may in fact be uniquely well positioned to tackle the issues plaguing PINNs. One exciting research direction is to achieve physics-informed learning by synergizing evolutionary algorithms with local gradient descent signals, thereby leveraging the best of both worlds to quickly arrive at physically accurate and generalizable PINN models. In what follows, we examine this and other promising algorithmic directions that we believe to warrant sustained investigation. These directions highlight key opportunities for the evolutionary optimization of PINNs (Evo-PINNs) to advance scientific machine learning.
@@ IV-A1 Neuroevolution Can Avoid Local Minima Traps
The operators of EAs are typically based on principles of natural selection. They do not require derivative computations of the loss function to be optimized and are therefore gradient-free. Importantly, as population-based methods, the implicit parallelism of EAs implies that they have lesser tendency to get trapped in a single local minimum [118, 49]. What is more, since their search is typically invariant to order-preserving transformations of the loss function [119], EAs are less susceptible to flat output functions with small (vanishing) gradients. For these reasons, they are deemed as potential alternatives to SGD, particularly in the context of scientific machine learning where the goal is to converge to physically precise solutions represented by global optima with zero PINN loss [48]. In addition to its global search capacity, an evolved population naturally provides an ensemble of neural networks that affords the possibility of accurate predictions with robust uncertainty quantification—as was shown through the application of a swarm-based evolutionary procedure to PINN optimization [120, 121]. The ability to provide uncertainty quantification and probabilistic forecasts is crucial for scientific modelling and control problems [122].
@@ IV-A2 Pareto Optimization Balances Physics and Data Losses
The existence of multiple loss terms—see Eqs. (2) and (3)—makes PINN training intrinsically a multi-objective optimization problem. Since the terms $\mathcal{L}_{data},\mathcal{L}_{pde},\mathcal{L}_{ic}$, and $\mathcal{L}_{bc}$ represent different physical quantities, their magnitudes may be incommensurable. In this case, the role of the scalarizing weights is not only to express preferences between the losses, but also to compensate for differences in their magnitudes. While a perfect physics solution would, in theory, simultaneously satisfy all PINN loss terms, the expressive capacity of a neural network is always limited by its (necessarily) finite size in practice. As argued in [89, 123], capacity limitations naturally give rise to conflicts and trade-offs between loss terms. This, in conjunction with the extreme ruggedness of the physics-informed learning objective, implies that a neural network, despite being a theoretical universal function approximator, is not guaranteed to find a solution that simultaneously minimizes all the terms [46]. When data-driven terms are also included, such as observational data or empirical measurements, true contradictions may arise as the data may be noisy, incomplete, or governed by partially unknown physics that differs from the modeled equations. In such cases, the multi-objective optimization must reconcile inherently conflicting objectives between data fitting and physics adherence. Identifying an appropriate aggregation of objectives to arrive at good (physically satisfactory) trade-off solution is also non-trivial [124] Moreover, the shape of the Pareto front formed by the multiple objectives is not known beforehand [125], with simple weighted-sum aggregations being insufficient to support all parts of non-convex fronts [126].
@@
Multi-objective EAs (MOEAs) are tailor-made to address these issues [127]. In particular, Pareto dominance-based variants of MOEAs [128] do not require the multiple loss terms to be scalarized, and hence are less sensitive to incommensurable objectives. Their population-based search mechanisms also allow for free-form convergence towards the entire Pareto front in a single run of the algorithm, irrespective of the front’s shape. These salient features have naturally begun to attract the attention of PINN researchers, although associated developments are still at a nascent stage. In [129], the non-dominated sorting genetic algorithm [130] was hybridized with a variant of SGD (namely, Adam [131]) for training the parameters of a PINN. The exploration capacity of the EA was shown to help overcome local minima and enable precise satisfaction of the initial and boundary conditions encoded in the physics-informed loss. Given the computational cost of a hybrid framework, a pure population-based multi-objective search algorithm was studied for PINN training in [121], leading to calibrated uncertainty quantification by examining the empirical distribution of the ensemble’s predictions.
@@
In addition to using MOEAs for PINN parameter training, there have been efforts towards evolving the model’s hyperparameters (including the scalarizing weights and activation functions). The method yields a Pareto set of hyperparameters where each setting offers an optimal trade-off between the data and physics losses. Practitioners can then decide post-hoc which setting they wish to pursue [132]. An analysis of such post-hoc decision-making reveals that solutions close to the convex bulge of the Pareto front typically lead to performant PINNs [124]. Evolutionary algorithmists will immediately notice a connection between this observation and the literature on knee-point driven MOEAs [133], which could serve as a valuable resource for hyperparameter tuning in PINNs.